OpenStudy (anonymous):

I'm having an issue solving my question. Can someone help? [15-20/2(2)]^-3

OpenStudy (anonymous):

\[\left[ 15-20\div2(2) \right]^{-3}\]

OpenStudy (anonymous):

I ended up getting 1/125 as my final answer. I'm not entirely sure if that is correct. I'm getting frusterated.

OpenStudy (anonymous):

it is hard to read, but i don't think that is right

OpenStudy (anonymous):

\(2\times 2=4, 20\div 4=5, 15-5=10, 10^{-3}=0.001\)

OpenStudy (anonymous):

But remember order of operations. 20/4 should come first and then 2x2. Then 1515-20

OpenStudy (helder_edwin):

i got -1/75